OpenStudy (anonymous):

I have another question. solve the equation: 3x-5=2x+7 So, do I combine like terms?

OpenStudy (anonymous):

you divide three on both sides

OpenStudy (mathteacher1729):

Are you trying to solve this algebraically? If so, you should subtract 2x from both sides: \[x-5=7\] Then add 5 to both sides \[x=12\] Test this to make sure it works: \[3(12)-5=2(12)+7\] \[36 -5=24+7\] \[31=31\] Yay! :D

OpenStudy (mathteacher1729):

I would strongly recommend graphing this to help understand that it's really the intersection of two lines. Let \[y_1=3x-5\] and \[y_2=2x+7\] You will see that they intersect at the exact (x,y) coordinate of (12,31) :)

OpenStudy (anonymous):

Yes, the answer in the book was x=12.. Thanks yet again!!

OpenStudy (mathteacher1729):

(on a TI calculator to find the intersection you may have to zoom out and then type 2nd > calc > intersect. ) Hope this helps!

OpenStudy (anonymous):

once I intersect I get x=12 and Y=31 on the TI